four oranges and five apples cost $3.56. Three oranges and four apples cost $2.76. find the cost of an apple and the cost of an orange.
Answer by jainenderkapoor(61)

Whenever this type of problem comes. Follow the simple steps mentioned below and get the answer.
Define two variables x and y (x representing cost of orange and y representing cost of apple)
Now if the cost of 1 orange is x, what is the cost of 2 oranges ??
YES it is 2x and COST of three oranges becomes 3x
So we can write the following two equations ---

4 x + 5 y = 3.56 ----- (1)
3 x + 4 y = 2.76 ----- (2)
Now our aim is to solve the two equations to get the value of x and y
We can use the either of the methods graphical methoid/ elimination method/ or equating the coefficients method
I am explaining the equating the coefficient method here.
We will be multiplying the first equation by `3`, which is the coefficient of x in the second equation, AND multiplying the equation 2 by `4`, which is the coefficient of x in first equation.
4 x + 5 y = 3.56 ----- (1) X 3
3 x + 4 y = 2.76 ----- (2) X 4

We get 12 x + 15 y = 10.68 ----- (3)
12 x + 16 y = 11.04 ----- (4)

On subtracting (3) from (4) we get
12 x + 16 y - 12 x - 15 y = 11.04 - 10.68
y = 0.36

Substituting the value of y in (1) or (2) we can get the value of x
(1) -- 4 x + 5 (.36) = 3.56
4 x + 1.8 = 3.56
4 x = 3.56 - 1.8 = 1.76
x = 1.76/4
x = 0.44
THIS IS HOW WE SOLVE THE QUESTION.
HOPE IT IS CLEAR TO YOU. IF YOUR NEED MORE HELP, I AM AVAILABLE FOR ONLINE TUTORING ALSO. YOU CAN CONTACT ME ON kapoorjai1@rediffmail.com